OpenStudy (anonymous):

Word Problem: How many liters of pure acid should a chemist mix with 4 L of a 10% acid solution to produce a solution that is 80% acid?

OpenStudy (anonymous):

0.8 = ((4 * 0.1) + A) / (4 * 0.9) + A)

OpenStudy (anonymous):

hmm lets try this. you have 4 liters that is 10% acid so it has 10% of 4 = .4 liters of acid in it. when you add say "x" liters you will have .4+x liters of acid and 4+x liters of liquid all together. you would like this 4+x liters to be 80% acid , that is you want .8(x+4) acid. so set these expressions equal to get \[.4+x=.8(x+4)\] and solve for x

OpenStudy (anonymous):

probably get exact same solution as above, but may be easier to see where the equation comes from. now solve. i would multiply by ten first to get rid of annoying decimals. \[4+10x=8(x+4)\] \[4+10x=8x+32\] \[2x=28\] \[x=14\] i think gives it, but you should check my algebra

OpenStudy (anonymous):

i think that's it. the first time i did the problem i got 28 though, but i think i she what i did wrong. Thanks so much.

OpenStudy (anonymous):

*see